OpenStudy (anonymous):

Stacey Steady pays $19.95 per month for her fitness center membership. If she works out twice a week for a year, how much does she pay for each visit?

OpenStudy (anonymous):

is there any answer choices

OpenStudy (anonymous):

19.95/2times*4 weeks= 19.95/8times/month

OpenStudy (anonymous):

Answer choices are: $1.92, $2.08, $2.30, and $2.49

OpenStudy (anonymous):

the answer is 2.49 because you take \[19.95/8=\]